The income expected to be received next year rises from $50,000 to $54,200. What will be consumption and saving if they were previously $21200 each


1
Expert's answer
2021-03-08T09:21:59-0500

Solution:

Let:

Y1 = Income in period 1 = 50,000

Y2 = Income in period 2 = 54,200

C1 = Consumption in period 1 = 21,200

S1 = Savings in period 1   = 21,200

C2 = Consumption in period 2?

S2 = Savings in period 2?

Therefore:

C2 = Y2 + S1 = 54,200 + 21,200 = $75,400

The consumption will be = $75,400

 

S2 = Y1 – C1 = 50,000 – 21,200 = $28,800

The savings will be = $28,800